OpenStudy (anonymous):

If y = 4x -2, find y when x = 4. A. 4 B. 6 C. 14

jimthompson5910 (jim_thompson5910):

Replace x with 4 and evaluate

jimthompson5910 (jim_thompson5910):

y = 4*4 - 2 y = ???

OpenStudy (anonymous):

14 :)

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

y= 4(4)-2 y= 16-2 so, y=14. :)